Calculate the pH of solutions containing the following:

a) 0.005 M HCl

b) one litre of 0.5M leucine buffer (pH 9.6) to which is added one litre of 0.1 M NaOH

c) 500 mL of 0.6 M H3PO4 plus 500 mL of 0.4 M KH2PO4 (pKa's for phosphate: 2.14,6.86, 12.4)

d) What is the total phosphate concentration in the buffer from part c)?
Answer #1

a) The pH of the strong acid is calculated:

pH = - log [HCl] = - log 0.005 = 2.30

c) By adding equal parts of each compound, the concentration drops by half, the pH is calculated using the equation of Henderson Hasselbach:

pH = pKa + log [Salt] / [Acid] = 2.14 + log (0.2 / 0.3) = 1.96

d) The total phosphate concentration is calculated:

[Phosphate] = [Acid] + [Salt] = 0.3 + 0.2 = 0.5 M

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.